# Josephson effect and self-trapping in helicoidal spin-orbit coupled Bose-Einstein condensates with optical lattices

### v1

@@ IV Population oscillation for non-zero zeeman splitting
Figure 8: Left panel: Variation of oscillation frequency with $\beta$ for (i) $\Delta=2$, $k_{n}=1.8$(dotted black) and (ii) $\Delta=0$, $k_{n}=1.8$(black dashed). Right panel: Variation of oscillation frequency of population imbalance with $k_{n}$ for (i) $\Delta=2$ and $\beta=1.6$ (black dotted) (ii) $\Delta=0$ and $\beta=1.5$(black dashed) and (iii) $\Delta=0$ and $\beta=1.6$(solid black). In both the panels we take, $G_{-}=1.05$ and $x_{0}=0.25$.
A comparative study on the oscillation dynamics of population imbalance for $\Delta=0$ and $\Delta\neq 0$ is shown in Fig.8. We see that frequency of oscillation is always larger in the case of $\Delta\neq 0$ than that for $\Delta=0$ (left panel). However, the difference in the frequency of oscillation between the two cases increases as $\beta$ increases for a fixed value of $k_{n}$ (right panel). The frequency of oscillation for a fixed value of $\beta$ shows similar trend in the cases of zero and non-zero Zeeman splitting. Particularly, it is larger at a smaller value of $k_{n}$ and it decreases gradually as $k_{n}$ increases for both $\Delta=0$ and $\Delta\neq 0$.
@@
Figure 10: Left panel:Variation of $<Z(t)>$ with $\Delta$ for $\varphi(0)=\pi/16$ and $\beta=1.6$. Middle panel: Variation of average value of atomic population imbalance ($<Z(t)>$) over a period with $\varphi(0)$ for $\Delta=4$ and $\beta=1.6$. Right panel: It gives variation of $<Z(t)>$ with $\beta$ for $\varphi(0)=\pi/16$ and $\Delta=2$.
The time evolution of density profile for non-zero Zeeman splitting clearly shows that the minimum values of atomic density in the two components are not equal Fig.9 (left panel and middle panel) and this inequality increases with the increase of $\Delta$. It indicates that self trapping occurs for $\Delta\neq 0$. To quantify the self trapping we calculate the value of population imbalance over a period ($<Z(t)>$) and show the results in Fig.10. We see that $<Z(t)>$ increases as $\Delta$ increases (left panel) for fixed values of $\varphi(0)$ and $\beta$ while $<Z(t)>$ decreases as $\varphi(0)$ increases (middle panel) if the values of $\Delta$ and $\beta$ are kept constant. For fixed values of $\Delta$ and $\varphi(0)$, however, $<Z(t)>$ increases with $\beta$ (right panel). Comparing the three cases we conclude that a larger self trapping can be obtained by changing the strength of helicoidal gauge potential.
@@
We consider a helicoidal spin-orbit coupled Bose-Einstein condensate embedded in optical lattices and studying the time evolution of population imbalance between two pseudo-spin 1/2 states. More specifically, we consider two important cases for the system, namely, the effects of (i) helicoidal gauge potential (ii) and optical lattices on population imbalance in a SOC-BEC for both zero and non-zero Zeeman splitting. In the case of zero Zeeman splitting, it is seen that the population imbalance oscillates in time about a point with equal population of the states (zero population imbalance). The amplitude of oscillation depends on the initial phase difference between the two states. In close analogy with the phase dependent current between two superconductor separated by a thin layer we refer the oscillation of atomic imbalance between spin-orbit coupled states, as Josephson-type oscillation. It shows that the frequency of JT oscillation increases as helicoidal gauge potential becomes weaker. In presence of optical lattices, the JT oscillation frequency for a fixed helicoidal gauge potential further increases and the amount of increment depends on the wavenumber of the lattice.
For non-zero Zeeman splitting the oscillation of population imbalance does not occur about its zero value. This results quantum mechanical self-tapping since the average value of population imbalance remains non-zero for a period of oscillation. The frequency of oscillation of population imbalance increases with the decrease of helicoidal gauge potential(HGP). We find that the self-trapping increases with the increase of HGP and Zeeman splitting. However, it decreases as initial phase difference increases. Thus we conclude that a pure JT oscillation can be observed in helicoidal SOC-BEC if the Zeeman splitting is zero.
@@ Acknowledgements
S. Sultana would like to thank ”West Bengal Higher Education Department” for providing Swami Vivekananda Merit Cum Means Scholarship with F. No. WBP201637916084.

### v3

@@ III Helicoidal gauge potentials and JT oscillation
Figure 3: Left panel:Variation of JT oscillation frequency with $s$. Here $G_{-}=1.05$. Right panel: Variation of $\delta\nu$ with $s$. In both the panels, we take $x_{0}=0.25$, $\Lambda=0.04$, $\alpha_{s}=4$ and $k_{n}=1.8$.
The effects of NOL on the JT oscillation in Halicoidal SOC-BEC is shown in Fig. 2. We see that the $Z(t)$ oscillates with time for different $\varphi(0)$ and the oscillations are quite similar to those observed in absence of NOLs. However, the frequency of oscillation ($\nu$) is larger than that found in the absence of NOL. This implies that the NOL causes acceleration in the dynamics of population imbalance.
The variation of $\nu$ for different lattice and SOC parameters for small oscillations can approximately be calculated from the relation given below.
$\nu=[A^{2}+A(\Lambda+B)]^{\frac{1}{2}}.$ (15)
In writing Eq. (15) we have used $\ddot{Z}+\nu^{2}Z=0$. Understandably, $\nu$ in Eq. (15) does not contain $\Delta$ since its dependence on $\Delta$ is nonlinear. We plot $\nu$ for unequal strength of spin-orbit coupling ($\alpha$) and helicoidal gauge potential ($\beta$). We see that $\nu$ decreases with the increase of $s(\alpha-\beta)$. This indicates that the spin-orbit coupling and helicoidal gauge potential oppose each other in changing population imbalance (left panel). As a result, $\nu$ decreases with the increase of $s$. The frequency of JT oscillation in presence and absence of NOL are different (right panel of Fig. 3) and the difference decreases as $s$ increases.
Figure 4: Density evolution of the first component $\rho_{1}=|\psi_{1}|^{2}$ (left panel), second component $\rho_{2}=|\psi_{2}|^{2}$ (middle panel) and total $\rho$= $\rho_{1}+\rho_{2}$ (right panel) for $\beta=1.6$ and $\varphi(0)=\pi/8$. Other parameters are kept same with those used in Fig. 3.
Variations of atomic densities with time in the two states of the system are shown in Fig. 4. It shows that the density of atomic population is changing between the two components (left and middle panel) periodically with time and the change of population imbalance is symmetric. This implies that maximum population in one component corresponds to minimum population of other component such that the total population remains constant with time (right panel).
IV Effects of helicoidal gauge potentials and optical lattices on self-trapping
In the case of JT oscillation, we have seen that population imbalance (Z(t)) oscillates with time about $Z=0$. Understandably, the oscillation of $Z$ about zero imbalance depends on the values of the different parameters of the system. We notice that there exists critical values of the parameters for observation of JT oscillation beyond which the oscillation does not occur about $Z=0$ and thus an asymmetry in the maximum population imbalance arises between the two states. Due to the asymmetry, the average value of population imbalance over one period becomes non-zero. This causes the so-called quantum mechanical self-trapping (QMST)[48]. In order to derive a condition for the occurrence of self-trapping, we recast Eqs. (12) and (13) in terms of Hamiltonian equation of motion, $\dot{Z}=-\partial H/\partial\varphi$ and $\dot{\varphi}=\partial H/\partial Z$ with the Hamiltonian given by
$H=\frac{(\Lambda+B)Z^{2}}{2}-A\sqrt{1-Z^{2}}\cos\varphi-\Delta Z.$ (16)
For fixed values of initial population imbalance and phase, the macroscopic self-trapping takes place if $H(Z(0),\varphi(0))$ exceeds the larger energy of the system[48]. Therefore, the critical condition for the observation of self-trapping can be obtained from Eq. (16) as
$\chi_{c}>\frac{2}{Z(0)^{2}}\big{[}1+\sqrt{1-Z(0)^{2}}\cos\varphi(0)+\frac{Z(0)\Delta}{A}\big{]}.$ (17)
Here, $\chi_{c}=(\Lambda+B)/A$. We see that, for a given value of $\Delta$, $Z(0)$ and $\varphi(0)$, the lattice and SOC parameters can be chosen properly for the observation of self-trapping. The inequality in Eq. (17) is termed as macroscopic quantum mechanical self-trapping (MQST) condition.
Variation of population imbalance with time for chosen values of different parameters satisfying the MQST condition is shown in Fig. 5. It shows that the population imbalance oscillates about a non-zero value and the amplitude of oscillation increases with the decrease of helicoidal gauge potential ($\beta$) (left panel). This implies that the self-trapping increases with the increase of halicoidal gauge potential. In order to quantify the amount of self-trapping, we calculate average value ($\langle Z(t)\rangle$) of population imbalance over a period for different values of $\beta$ and see that $\langle Z(t)\rangle$ increases nonlinearly as $\beta$ increases(right panel).
Figure 5: Left panel: Variation of population imbalance with $t$ for $\beta=3.2$ (solid blue) and $\beta=4$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $\beta$. In both the panels, we take $\Lambda=1.817$, $\alpha_{s}=4$, $x_{0}=0.25$, $G_{-}=2.05$, $k_{n}=1.8$ and $\Delta=0$.
We examine the effects of interaction parameter $\Lambda$ through the variation of population imbalance with time by taking $\Lambda=0.665$ (solid blue) and $\Lambda=1.26$ (solid black) in Fig.6. The value of $\Lambda$ are so chosen that the binary condensates remain miscible. We see that the amplitude of oscillation decreases with the increase of $\Lambda$. This indicates that the self-trapping is influenced by the relative value of inter- and intra-atomic interactions. We calculate $\langle Z(t)\rangle$ for different values of $\Lambda$ and see that the self-trapping grows linearly up to $\Lambda\approx 1.2$.
@@ IV Effects of helicoidal gauge potentials and optical lattices on self-trapping
In order to find the effects of nonlinear optical lattices on the self-trapping, we plot $Z(t)$ for two different lattice strengths, namely, $G_{-}=2.05$ (solid blue)and $G_{-}=3.05$ (solid black) in Fig 7. We know that the nonlinear optical lattice influences the phase velocity of $Z(t)$ oscillation and thus plays an important role in changing the frequency of the oscillation [20]. The variation of frequency of oscillation ($\nu$) for different values of NOLs’ strengths is clearly reflected in the left panel of Fig. 8. More specifically, $\nu$ increases as the difference of lattice strengths ($G_{-}$) increases. The population imbalance in this case also oscillates about $Z(t)\neq 0$ and the amplitude of oscillation increases with the decrease of $G_{-}$. We calculate $\langle Z(t)\rangle$ for different values of $G_{-}$. The result displayed in Fig. 7 (right panel) shows that the self-trapping first increases and then attains a saturated value if $G_{-}$ further increases. We have checked that the self-trapping decreases if we take as $\Delta\neq 0$ for particular values of lattice and SOC parameters (Fig. 8).
Figure 7: Left panel: Variation of population imbalance with $t$ for $G_{-}=2.05$ (solid blue)and $G_{-}=3.05$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $G_{-}$. In both the panels we take $\beta=2$, $\alpha_{s}=2$, $x_{0}=0.25$, $k_{n}=1.8$, $\Lambda=0.681$ and $\Delta=0$.
Figure 8: Left panel: Variation of population imbalance with $t$ for $\Delta=0$ (solid blue) and $\Delta=1.2$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $\Delta$. In both the panels we take $\beta=5$, $\alpha_{s}=0.2$, $x_{0}=0.25$, $k_{n}=0.8$, $\Lambda=1.81$ and $G_{-}=2.05$.
Figure 9: Density evolution of the first component $\rho_{1}=\lvert\psi_{1}\lvert^{2}$ (left panel), second component $\rho_{2}=\lvert\psi_{2}\lvert^{2}$ (middle panel) and total $\rho$= $\rho_{1}+\rho_{2}$ (right panel) for $\alpha_{s}=4$, $x_{0}=0.25$, $\beta=3.2$ and $\Delta=0$.
The variation of atomic density($\rho_{j}$=$\lvert\psi_{j}\lvert^{2}$) with time during quantum mechanical self-trapping is shown in Fig. 9. It clearly shows that the maximum density of one component corresponds to the minimum of others. However, the values of maxima or minima of atomic densities of two components are not equal and thus result MQST.
@@
We consider a helicoidal spin-orbit coupled Bose-Einstein condensate embedded in optical lattices and studying the time evolution of population imbalance between two pseudo-spin 1/2 states. More specifically, we consider two important cases for the system, namely, the effects of (i) helicoidal gauge potential and (ii) nonlinear optical lattices on population imbalance in a spin-orbit coupled Bose-Einstein condensate. It is seen that the population imbalance oscillates in time about a point with zero population imbalance. The amplitude of oscillation depends on the initial phase difference between the two states. In close analogy with the phase dependent current between two superconductor separated by a thin layer we refer the oscillation of atomic imbalance between spin-orbit coupled states as Josephson-type (JT) oscillation. We have seen that the frequency of oscillation increases as helicoidal gauge potential becomes weaker. In presence of nonlinear optical lattices, the frequency of Josephson-type oscillation for a fixed helicoidal gauge potential further increases.
We have examined that the oscillation of population imbalance does not occur about zero imbalance if the parameters of the system do not satisfy the so-called macroscopic quantum mechanical self-trapping condition. We quantify the self-trapping by calculating average value of population imbalance for different values of (i) helicoidal gauge potential, (ii) nonlinear lattice parameters and (iii) inter and itra-atomic interaction strengths. We find that, for a particular value of helicoidal gauge potential, self-trapping grows appreciably with the change of atomic interaction and lattice parameters. However, we observe saturation in the self-trapping for higher values of atomic interactions.
@@ Acknowledgements
S. Sultana would like to thank ‘West Bengal Higher Education Department’ for providing Swami Vivekananda Merit Cum Means Scholarship with F. No. WBP201637916084.